OpenStudy (anonymous):

Find the solution set. Separate the two calues with a comma: x^2+6x+8=0

OpenStudy (amistre64):

x^2 + (m+n)x + (mn) = 0 is the setup we need to find m and n

OpenStudy (amistre64):

m and n are factors of 8; 1(8); 2(4) m+n are the two factors of 8 that add up to 6 1+8 = 9 2+4 = 6

OpenStudy (anonymous):

use the quadratic formula: \[x = -3 \pm \sqrt{9-8} = -3 \pm 1 =(-4,-2)\]

OpenStudy (amistre64):

(x+m)(x+n) = 0 (x+4)(x+2) = 0 when x+4 = 0 or x+2 = 0

OpenStudy (anonymous):

Thanks XaLeX